OpenStudy (anonymous):

below are two equations . solve each equation and compare the two solutions . choose the statement that is true about each solution

OpenStudy (jiteshmeghwal9):

Please post the equations @tiff1998g :)

OpenStudy (anonymous):

equation 1 : 8- 5x = 4x - 1 equation 2 : 4x + 4 = 10 - 2x

OpenStudy (anonymous):

The equations are equivalent.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Are you able to solve the equations and see that they are equivalent?